Light with wavelength 582 nm goes through a single slit of width 0.026 mm and displays a diffraction pattern on a screen 4.2 m away. 12) What is the width of the central bright fringe? a) 0.09 m b) 0.19 m c) 0.29 m d) 0.39 m e) 0.49 m
Answer #1

Given light wavelength lambda = 582 nm , slit width d = 0.026 mm

sepration of slit and screen is D = 4.2 m

we know that the central fringe is maximum and with more intensity

12)

the width of the central fringe is W = LAmbda*D/d

W = 582*10^-9*4.2/(0.026*10^-3) m

W = 0.0940 m

Answer is option (a)

13)

3rd dark fringe angle theta = ?

d sin theta = (m*lambda )

0.026*10^-3 sin theta = 3*582*10^-9

solving for theta , theta = 3.85 degrees = 3.9 degrees

answer is opiton (b)

14 ) if frequency increases the fringes would get closer together

answer is option (b)